OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL!!! Solve the system of equations algebraically. Show all your steps. y=x^2+2x y=3x+20

OpenStudy (anonymous):

@satellite73 @chris215 @Gabylovesyou

OpenStudy (anonymous):

\[\text{Solve}\left[x^2+2 x=3 x+20\right] \]\[{{x = -4}, {x = 5}}\]

OpenStudy (anonymous):

Use the x values above to solve for the corresponding y values.